The initial prssure of COCl2 is 1000 torr.The total pressure of the system becomes 1500 torr.when the equilb COCl2(g)=CO(g) + Cl2(g).(reversible) is attained at constant temp.The value of Kp ?

Grade:12

2 Answers

Naveen Kumar
9 years ago
The given reaction is :
CoCl2(g).......................................=.......CO(g).....+.....Cl2(g)
1000 torr..................................................0...............0.....................(initially)
1000-p.....................................................p...............p....
let p is the pressure of CO as well as of CL2.n so same partial pressure of COCl2 will be decreased
So total pressure at equilibrium=1000-p+p+p=1000+p=1500
Now calculate p .
Writing, Kp=pCO*pCl2/pCOCl2
=p*p/(1000-p)
now we can get Kp

Ramreddy IIT Patna
9 years ago
COCl2 -------> CO + Cl2
1000 torr (initial)
1000(1-x) 1000x 1000x
Total pressure= 1000(1+x)
= 1500
x= 0.5 or (50% dissociation)
Kp= (pCl2)(pCO)/(pCOCl2)
= 500 X 500 /500
you will get the answer.
….….….….….….….….….….….….….….….….….….….….…....